OpenStudy (anonymous):

find the slope of the line tangent to the graph of the function g(x)=9-x^2 at the point (4,-7)

OpenStudy (raden):

slope = g'(x = 4) find g'(x), first

OpenStudy (anonymous):

the derivative of 4 is 0 @RadEn

OpenStudy (raden):

g(x) = 9 - x^2 g'(x) = -2x then slope = g'(x=4) = -2(4) = -8

OpenStudy (anonymous):

thank you!

OpenStudy (raden):

welcome

OpenStudy (anonymous):

ok @RadEn the next question is find the slope of the line tangent to the graph of the function f(x)=3-7x at the point (-1,10). would it be -7?

OpenStudy (raden):

you are right :)

OpenStudy (anonymous):

thanks!

OpenStudy (raden):

you're welcome